OpenStudy (anonymous):

I need major help with my math homework tonight. It is on Graphing Linear Equations. The directions say "Graph each equation by making a table" and My math teacher says I need to get the Y alone on this problem and then graph it using a table. 10. -3+2y=-5 And Here is an example problem she helped me with earlier. 8. 2x-5y=1 X intercept: 2x-5(0)=1 2x/2=1/2 x=1/2 (1/2,0) Y intercept: 2(0)+-5y=1 -5y/-5=1/-5 y=-5 (0,-1/5) So i graphed (1/2,0) and (0,-1/5) and got a liner line. She said during the lesson that the X&Y values always equals 0.

OpenStudy (anonymous):

What exactly is the problem you face?

OpenStudy (anonymous):

Solve the equation for y

OpenStudy (anonymous):

10. -3+2y=-5

OpenStudy (anonymous):

youll get y=-1

OpenStudy (anonymous):

-3+2y=-5 -3+3+2y=-5+3 2y=-2 2y/2=-2/2 y=-1